# Data-driven Software-based Power Estimation for Embedded Devices

### v1

@@ Abstract
Energy measurement of computer devices, which are widely used in the Internet of Things (IoT), is an important yet challenging task. Most of these IoT devices lack ready-to-use hardware or software for power measurement. A cost-effective solution is to use low-end consumer-grade power meters. However, these low-end power meters cannot provide accurate instantaneous power measurements. In this paper, we propose an easy-to-use approach to derive an instantaneous software-based energy estimation model with only low-end power meters based on data-driven analysis through machine learning. Our solution is demonstrated with a Jetson Nano board and Ruideng UM25C USB power meter. Various machine learning methods combined with our smart data collection method and physical measurement are explored. Benchmarks were used to evaluate the derived software-power model for the Jetson Nano board and Raspberry Pi. The results show that 92% accuracy can be achieved compared to the long-duration measurement. A kernel module that can collect running traces of utilization and frequencies needed is developed, together with the power model derived, for power prediction for programs running in real environment.
@@
Power Estimation, Machine Learning
@@ I Introduction
  • Energy consumption is an important indicator of system performance [1] [2] [3]. In the realms of head-mounted AR [4] and edge computing [5], energy efficiency directly influences the operational capabilities and sustainability.
  • Some energy-aware system design methodology makes control based on energy consumption feedback [6] [7] [8].
@@
Unfortunately, as of today, many popular embedded hardware (for example, Raspberry Pi and Nvidia Jetson Board) [12] [18] [19] that is widely used in IoT and Mobile Edge Computing (MEC) System, lacks a software-level approach for energy measurement. In this case, users who do not have access to high-end equipment can only measure power consumption based on consumer-grade meters [20] [21] [22]. We next discuss the limitations of using consumer-grade meters.
@@ I-A Limitations of using consumer-grade power meters for energy measurement
Consumer-grade power meters often do not provide data API. To integrate power measurement into energy-saving policy design, a practical solution will be using the data interface provided by the meter (similar to the sysfs interface in Linux) for programs to access power meter measurement. Unfortunately, we did not find a consumer-grade energy meter that offers a data API. In most cases, the power meter provides a GUI interface for the user to read data from the hardware or software connected to the meter via Bluetooth [23]. The lack of a low-overhead and low-latency data access scheme means that making a policy design approach feasible based on energy feedback for control is challenging with consumer-grade power meters.
@@
Due to hardware limitations, consumer-grade meters record voltage/current changes in a coarse-grained view, usually in seconds. This means it is difficult to accurately measure energy consumption over short periods of time. Next we illustrate this point using an example of measuring the energy consumption of a CPU-intensive program with an execution time of around 1 second. We applied two methods. One is based on short-duration measurement, where we simply record the energy consumption during the 1-second execution. Another one is based on long-duration measurement, where we run the workload repetitively in a loop for three minutes and obtain the average energy consumption for 1-second execution as the total 2-minute energy consumption divided by the number of executions. The test is shown in Fig. 1. The variance of measurements over short periods of time is much larger than over long periods of time. This illustrates that short-duration measurement is not reliable and only measurements over a long period of time can provide an accurate, consistent result.
Because of the above-discussed limitations, energy measurements in our previous works [24] [25], were very labor intensive.
I-B Proposed approach
In this paper, we propose a data-driven method to derive software-based power estimate model using only consumer-grade meter. The contributions of the paper can be summarized as follows.
  • For users or researchers with devices that do not have software-based energy measurement support and have access only to consumer-grade meters, our method can help them construct a software-based energy consumption estimation model for their devices.
  • Our estimation model allows short-duration estimation, overcoming the limitation of the consumer-grade power meters.
@@ I-C Related works
Unlike design solutions that are based on mathematical models for offline analysis[26] [27] [28] [29] [30] [31] [32] [33] [34], adaptive system design and measurement methods based on machine learning [24] [25] [35] [36] [37] [38] [39] [40] [41] [42] [43] [44] have received attention in recent years. Accurate power modeling and measurement have been widely explored in the literature, with various approaches proposed for different types of processors and systems. Below, we review some key contributions and compare them with our proposed method.
@@ II Data-driven Power Estimation with Consumer-grade Meters
The hardware connection for the approach is shown in Figure 2. The power meter is connected to the underlying device to be studied. Note that workload will run on the device for data collection to construct energy model for the device. A PC is used to receive the Bluetooth data sent by the power meter, which is energy consumption data. In addition, the parameters (frequency and utilization of the device) that are used to build the model, are recorded when the workload runs on the device. These data are integrated together with the energy consumption data for energy model construction. A data collection module is required to retrieve the necessary parameters like the frequency and utilization data for using the model.
@@ II-B Derive Power Model from Collected Data
After obtaining the training data, various methods can be explored to derive a model to predict the power for a given utilization and frequency. The models that we explored include regression models, decision tree models, and neural network models. Note that the types of models to be used are not limited to the above. The details of deriving models for a Jeston Nano board will be described in section III. Here we show the regression model as an example.
$Power=$ $k_{0}\cdot utilization^{2}+k_{1}\cdot frequency\cdot utilization$ (1)
$+k_{2}\cdot utilization+b$
Different regression models were explored for the Jeston nano board. Among them, the best results were obtained from the per-frequency regression model. In this model, at each frequency the device supports, there is a polynomial expression consisting of utilization, as shown in Equation 1.
At each frequency, there are corresponding values for $k_{0}$, $k_{1}$, $k_{2}$ and $b$. The reason for choosing such a regression model is that the change in frequency causes a jump of the power, which will be explained in detail in Section III-C1.
@@ II-C How to Use the Software-based Power Model?
To predict the power for programs running with multiple frequencies, it is necessary to obtain multiple frequencies for the running trace, the utilization for each frequency and the duration of the frequency. With the utilization of each frequency segment, the power over this frequency period can be obtained based on our model.
@@
To get the duration and utilization information for a running traces, we built a kernel module called DataTracker module that can directly track all the frequencies used together with their duration and the utilization of each core at each frequency change. The detailed information of this module will be explained in Section IV-A.
@@ III-A4 Training Data Illustration
The embedded device, Nvidia Jetson Nano Board 2GB (JTN), supports a range of frequencies from 0.102 GHz to 1.479 GHz. The number of CPU cores is 4. Training data collection is done based on the algorithm explained in Section II-A. The training data are shown in Fig. 3. This is a 3D figure with the x, y, and z axes representing frequency, utilization, and power, respectively.
@@
It can be seen that as the frequency and utilization increase, so does the power. When the frequency is constant, the utilization and power are smoothly related. When the utilization is roughly within a range, there is a jump in power as the frequency changes. This is obvious in the single-threaded benchmark. The first three frequencies correspond to a power level, and the latter frequency corresponds to another power level. We inferred that the reason for this jump is the change in voltage. Comparing the single-threaded data with the multi-threaded data, we can see that the multi-threaded data has a steeper trend.
@@ III-B Testing Data Collection
We chose to select benchmarks from Mibench [56] and Sysbench [57] as the testing data. These two test sets were chosen because they are both commonly used, lightweight and easy to measure. Mibench contains different types of benchmarks, most of them are single-threaded, while the benchmarks in Sysbench are multi-threaded.
We selected 10 benchmarks from Mibench, which are $basicmath$, $bitcount$, $jpeg$, $typeset$, $dijkstra$, $particia$, $stringsearch$, $blowfish$, $sha$ and $crc32$. These 10 benchmarks cover all types in Mibench.
@@
We set the CPU frequency to each available level when running each Mibench (or Sysbench) benchmark. We then obtain the utilization and power for each running case. The utilization is obtained by reading from the $/proc$ file system, and the power is obtained by a long-time measurement with the power meter. At each frequency, each benchmark is run multiple times for 3 minutes. The purpose of this is to reduce errors introduced by single or short-time measurement. The method of reading energy consumption is the same as in Section II-A. Fig. 4 shows the measurement results of the Mibench benchmarks.
@@
Figure 4: Measurement results of 9 benchmarks in Mibench
@@
The one remaining benchmark in Mibench, $stringsearch$, has a different trend from these nine benchmarks, as shown in Fig. 5. The other benchmark is $bitcount$ of the other 9, selected to be used as a comparison. It can be seen from the figure that the power of $bitcount$ is shown as the previous illustration where it has a jump. However, the power of $stringsearch$ show a different trend. Its power consumption increases progressively with increasing frequency. The reason is that the other benchmarks show little change in utilization at all frequencies. However, $stringsearch$ has higher utilization at lower frequencies than the other benchmarks. At high frequencies, its utilization is reduced. Therefore no jump in energy consumption of this benchmark.
@@ III-C Model Derivation and Evaluation
We use the data collected (illustrated in Section III-A4) to build the models and the test data from Section III-B to evaluate the accuracy of the models. The metrics used are mean squared error (MSE), mean absolute error (MAE), and $R^{2}$ score [58].
Mean Squared Error (MSE) is the average of the squares of the errors—that is, the average squared difference between the estimated values and the actual value. Mean Absolute Error (MAE) is the average of the absolute errors—the average absolute difference between the estimated values and the actual value. The $R^{2}$ score indicates the degree of explanation of the dependent variable by the independent variable. It ranges from 0 to 1. The higher the value of this indicator, the higher the degree of prediction accuracy.
@@ III-C1 Polynomial Regression Model
We constructed four linear models based on the training data. The first model we choose is the simplest polynomial model. We call Model 1 the Simple Regression Model. Frequency and utilization as two parameters are selected in the model as cubic and primary, respectively. This is because, in the CMOS circuit model, frequency and power consumption are cubic. The second model is optimized based on Model 1. It added more terms with different exponents, with the highest order being 3. The reason for this modification is simply the intuitive idea that adding more terms will improve the accuracy of the model. We call Model 2 the Multi-Term Regression Model. Considering the frequency/voltage levels, we further divide Model 2 into two different equations corresponding to 2 frequency/voltage levels. It uses one set of coefficients when the frequency is in the first 3 levels, and another set of coefficients when the frequency is in the later levels. We call it the Multi-Frequency Regression Model. We improve Model 3 further and introduce Model 4, called the Per-Frequency Regression Model, which fits data to individual frequency.
All results of four models based on testing data can be seen from Table I.
TABLE I: The accuracy of regression models
@@
Simple Regression Model 0.2397 0.4184 -0.0242
Multi-Term Regression Model 0.0272 0.1365 0.8839
Multi-Frequency Regression Model 0.0270 0.1344 0.8848
Per-Frequency Regression Model 0.0182 0.1040 0.9221

### v2

@@ Abstract
Energy measurement of computer devices, which are widely used in the Internet of Things (IoT), is an important yet challenging task. Most of these IoT devices lack ready-to-use hardware or software for power measurement. In this paper, we propose an easy-to-use approach to derive a software-based energy estimation model with external low-end power meters based on data-driven analysis. Our solution is demonstrated with a Jetson Nano board and Ruideng UM25C USB power meter. Various machine learning methods combined with our smart data collection & profiling method and physical measurement are explored. Periodic Long-duration measurements are utilized in the experiments to derive and validate power models, allowing more accurate power readings from the low-end power meter. Benchmarks were used to evaluate the derived software-power model for the Jetson Nano board and Raspberry Pi. The results show that 92% accuracy can be achieved by the software-based power estimation compared to measurement. A kernel module that can collect running traces of utilization and frequencies needed is developed, together with the power model derived, for power prediction for programs running in a real environment. Our cost-effective method facilitates accurate instantaneous power estimation, which low-end power meters cannot directly provide.
@@
Power Estimation, Smart Profiling, Machine Learning
@@ I Introduction
  • Energy consumption is an important indicator of system performance [1] [2] [3]. In the realms of head-mounted AR [4] and edge computing [5], energy efficiency directly influences operational capabilities and sustainability.
  • Some energy-aware system design methodology makes control based on energy consumption feedback [6] [7].
@@
Unfortunately, as of today, many popular embedded hardware that is widely used in IoT and Mobile Edge Computing (MEC) System, lacks a software-level approach for energy measurement. In this case, users who do not have access to high-end equipment can only measure power consumption based on consumer-grade meters [17] [18]. We next discuss the limitations of using consumer-grade meters.
@@ I-A Limitations of using consumer-grade power meters for energy measurement
Consumer-grade power meters often do not provide data API. To integrate power measurement into energy-saving policy design, a practical solution will be using the data interface provided by the meter (similar to the sysfs interface in Linux) for programs to access power meter measurement. Unfortunately, we did not find a consumer-grade energy meter that offers a data API. In most cases, the power meter provides a GUI interface for the user to read data from the hardware or software connected to the meter via Bluetooth [19]. The lack of a low-overhead and low-latency data access scheme means that designing low-power management methods based on energy feedback for control is challenging with consumer-grade power meters.
@@
Due to hardware limitations, consumer-grade meters record voltage/current changes in a coarse-grained view, usually in seconds. This means it is difficult to accurately measure energy consumption over short periods of time. Next we illustrate this point using an example of measuring the energy consumption of a CPU-intensive program with an execution time of around 1 second. We applied two methods. One is based on short-duration measurement, where we simply record the energy consumption during the 1-second execution. Another one is based on long-duration measurement, where we run the workload repetitively in a loop for three minutes and obtain the average energy consumption for 1-second execution as the total 3-minute energy consumption divided by the number of executions. The test is shown in Fig. 1. The variance of measurements over short periods of time is much larger than over long periods of time. This illustrates that short-duration measurement is not reliable and only measurements over a long period of time can provide an accurate, consistent result.
Because of the above-discussed limitations, energy measurements in our previous works [20] [21], were very labor intensive. The work proposed in this paper overcomes the above limitations.
I-B Proposed approach: data-driven power estimation
In this paper, we propose a data-driven method to derive software-based power estimate models using affordable external consumer-grade meters, and provide a profiling tool to estimate the power of programs. The contributions of the paper can be summarized as follows.
  • We propose a method to construct a software-based energy consumption estimation model for small devices.
@@ I-B Proposed approach: data-driven power estimation
  • We also provide a platform-independent kernel module that can be used to collect running traces of utilization and frequencies needed for power prediction for programs running in Linux environment.
  • The approach can help users estimate energy on devices lacking integrated energy measurement software, especially when they only have access to affordable consumer-grade meters. Our approach takes into account short-duration estimation error, and overcomes the limitation of the consumer-grade power meters.
@@ I-C Related works
Edge computing on small devices has become a crucial component of today’s intelligent society and has attracted significant research interest [3] [5] [14] [16] [17] [21] [18] [20] [22] [23] [24] [25] [26] [27]. In system design, unlike design solutions that are based on mathematical models for offline analysis[22] [28][29], adaptive system design and measurement methods [20] [21] [24] [25] [26] [30] [31] [32] [33] [34] [27] [35] [36] [37] [38] have received attention in recent years. Accurate power modeling and measurement have been widely explored in the literature, with various approaches proposed for different types of processors and systems. Below, we review some key contributions and compare them with our proposed method.
@@
Table I and Table II summarize the key parameters used in the power estimation models and energy acquisition methods for different approaches. Note that PMC stands for Performance Monitoring Counter.
TABLE I: The Key Parameters for Power Estimation
Types Works
Register Transfer Level (RTL) Model [42], [43]
Micro-Architecture Model + Code Trace [48] + [49]
Instruction Level Energy Model [50]
PMC: Executed instructions, branch … [41]
PMC: Instruction per Cycle, Cache Miss Rate, … [39]
PMC: Select events for processor-specific configuration [40]
Frequency, Utilization Our Work
TABLE II: Energy Acquisition Methods
Types Works
High Precision Instrument [39]
Custom Board [50] [41]
Build-in Energy Sensor [40]
Consumer Grade Meter Our Work
Simulator [44] [45] [46] [47]
@@ II Data-driven Power Estimation with Consumer-grade Meters
The hardware connection for the approach is shown in Figure 2. The power meter is connected to the underlying device to be studied. Note that workload will run on the device for data collection to construct energy model for the device. A PC is used to receive the Bluetooth data sent by the power meter, which is energy consumption data. In addition, the parameters (frequency and utilization of the device) that are used to build the model, are recorded when the workload runs on the device. These data are integrated with the energy consumption data for energy-model construction. A data collection module is required to retrieve the necessary parameters like the frequency and utilization data for using the model.
@@ II-B Derive Power Model from Collected Data
After obtaining the training data, various methods can be explored to derive a model to predict the power for a given utilization and frequency. The models that we explored include regression models, decision tree models, and neural network models. Note that the types of models to be used are not limited to the above. The details of deriving models for a Jeston Nano board will be described in section III. Here we show the regression model as an example, where $util$ denotes utilization and $freq$ denotes frequency.
$Power=k_{0}\cdot util^{2}+k_{1}\cdot freq\cdot util+k_{2}\cdot util+b$ (1)
Different regression models were explored for the Jeston nano board. Among them, the best results were obtained from the per-frequency regression model. In this model, at each frequency the device supports, there is a polynomial expression consisting of utilization, as shown in Equation 1. At each frequency $i$, the corresponding coefficients are $k_{0}^{i}$, $k_{1}^{i}$, $k_{2}^{i}$ and $b^{i}$. The reason for choosing such a regression model is that the change in frequency could cause a jump in power. More details can be found in Section III-C1.
@@ II-C How to Use the Software-based Power Model?
To predict the power of a program running with multiple frequencies, it is necessary to obtain all the frequencies used running the program, the utilization for each frequency, and the duration of the frequency. With the utilization of each frequency segment, the power over this frequency period can be obtained based on our model.
@@
To get the duration and utilization information for running traces, we built a kernel module called DataTracker module that can directly track all the frequencies used together with their duration and the utilization of each core at each frequency change. The detailed information of this module will be explained in Section IV-A.
@@ III-A4 Training Data Illustration
The embedded device, Nvidia Jetson Nano Board 2GB (JTN), supports a range of frequencies from 0.102 GHz to 1.479 GHz. The supported utilisation range is 0% to 100%. The number of CPU cores is 4. Training data collection is done based on the algorithm explained in Section II-A. The training data are shown in Fig. 3. This is a 3D figure with the x, y, and z axes representing frequency, utilization, and power, respectively.
@@
It can be seen that as the frequency and utilization increase, so does the power. When the frequency is constant, the utilization and power are smoothly related. When the utilization is roughly within a range, there is a jump in power as the frequency changes. This is obvious in the single-threaded benchmark. The first three frequencies correspond to a power level, and the latter frequency corresponds to another power level. We inferred that the reason for this jump is the change in voltage. Comparing the single-threaded data with the multi-threaded data, we can see that the multi-threaded data have a steeper trend.
@@ III-B Testing Data Collection
We chose to select benchmarks (see Table III for details) from Mibench [52] and Sysbench [53] as the testing data. These two test sets were chosen because they are both commonly used, lightweight and easy to measure. Mibench contains different types of benchmarks, most of them are single-threaded, while the benchmarks in Sysbench are multi-threaded. The 10 benchmarks (Table III) selected from Mibench cover all the types in Mibench, aiming to mitigate bias.
@@
TABLE III: List of testing benchmark programs
Suite Benchmark Description
Mibench basicmath Basic mathematical calculations
bitcount Bit manipulation and counting
jpeg JPEG image compression and decompression
typeset Typesetting operations
dijkstra Dijkstra’s shortest path algorithm
patricia Patricia trie data structure algorithm
stringsearch String searching algorithms
blowfish Blowfish encryption algorithm
sha Secure Hash Algorithm (SHA) computation
crc32 CRC32 checksum computation
Sysbench cpu CPU-intensive computational tasks
memory Memory-related operations
fileio File read and write operations
We set the CPU frequency to each available level when running each Mibench (or Sysbench) benchmark. We then obtain the utilization and the measured power for each running case. The utilization is obtained by reading from the $/proc$ file system, and the power is obtained by a long-time measurement with the power meter. At each frequency, each benchmark is run multiple times for 3 minutes. The purpose of this is to reduce errors introduced by single or short-time measurements. The method of reading energy consumption is the same as in Section II-A. Fig. 4 shows the measurement results of the Mibench benchmarks.
@@
Figure 4: Measurement results of 10 benchmarks in Mibench
@@
The $stringsearch$ benchmark in Mibench has a different trend than the other nine benchmarks. In Fig. 5, one of the nine benchmarks, $bitcount$, is used to show the difference with $stringsearch$. It can be seen from the figure that the power of $bitcount$ has a jump (a similar trend can be found in the 9 benchmarks other than $stringsearch$ in Fig. 4). On the other hand, the power of $stringsearch$ increases progressively with increasing frequency. We can observe that the nine benchmarks show little change in utilization at all frequencies while $stringsearch$ has higher utilization at lower frequencies than the other benchmarks. At high frequencies, $stringsearch$’s utilization is reduced. Therefore no jump in energy consumption of this benchmark.
@@ III-C Model Derivation and Evaluation
We use the data collected (illustrated in Section III-A4) to build the models. We have explored linear regression-based, decision-tree-based and neural-networked-based software models for power estimation, which will be described in detail in the later part of this section.
We use the benchmarks in Table III to evaluate the prediction accuracy of the models. The metrics used are mean squared error (MSE), mean absolute error (MAE), and $R^{2}$ score [54]. Mean Squared Error (MSE) is the average of the squares of the errors—that is, the average squared difference between the estimated values and the actual value. Mean Absolute Error (MAE) is the average of the absolute errors—the average absolute difference between the estimated values and the actual value. The $R^{2}$ score indicates the degree of explanation of the dependent variable by the independent variable. It ranges from 0 to 1. The higher the value of this indicator, the higher the degree of prediction accuracy.
Algorithm 2 Model Evaluation for Model $M$
1:Set $governor$ to userspace
2:Initialize $MSE\leftarrow 0$
3:Initialize $MAE\leftarrow 0$
4:Initialize $n\leftarrow 0$
5:Initialize $SumP\leftarrow 0$
6:Initialize empty lists $P\_list$ and $P^{M}\_list$
7:for each benchmark program $i$ do
8:     for each frequency $j$ that the device supports do
9:         Run benchmark $i$ at frequency $j$
10:         Measure utilization $u_{ij}$ and actual power $P_{ij}$
11:         Calculate predicted power $P^{M}_{ij}$ using model $M$ with $u_{ij}$ and frequency $j$
12:         $MSE\leftarrow MSE+(P_{ij}-P^{M}_{ij})^{2}$
13:         $MAE\leftarrow MAE+|P_{ij}-P^{M}_{ij}|$
14:         Append $P_{ij}$ to $P\_list$
15:         Append $P^{M}_{ij}$ to $P^{M}\_list$
16:         $SumP\leftarrow SumP+P_{ij}$
17:         $n\leftarrow n+1$
18:     end for
19:end for
20:Compute mean actual power $\overline{P}=\dfrac{SumP}{n}$
21:Compute $MSE\leftarrow\dfrac{MSE}{n}$
22:Compute $MAE\leftarrow\dfrac{MAE}{n}$
23:Initialize $SSR\leftarrow 0$triangleright Sum of Squares of Residuals
24:Initialize $SST\leftarrow 0$triangleright Total Sum of Squares
25:for $k\leftarrow 1$ to $n$ do
26:     $SSR\leftarrow SSR+(P\_list[k]-P^{M}\_list[k])^{2}$
27:     $SST\leftarrow SST+(P\_list[k]-\overline{P})^{2}$
28:end for
29:Compute coefficient of determination $R^{2}=1-\dfrac{SSR}{SST}$
How to evaluate a given power estimation Model $M$ with the above metrics is shown in Algorithm 2. To obtain the measured values for comparison with the predicted values, we configured each benchmark to run at various frequency settings and measured the corresponding utilization and power consumption. Specifically, we recorded each measurement in the format of benchmark identifier, utilization, frequency, and power. Given that there are multiple frequency settings (referred to as frequency steps) and several different benchmarks, this process yielded a total number of data points ($n$ in Algorithm 2) equal to the product of the number of frequencies and benchmarks. After collecting the actual power consumption data, we applied the models discussed in the following sections to predict the power values. We then compared the actual power data with the predicted power data to evaluate the performance indicators.
Next, we describe the software power estimation models explored on the Jetson device.
@@ III-C1 Polynomial Regression Model
We constructed four linear models based on the training data. The first model (Model 1) we choose is the simplest polynomial model. We call Model 1 the Simple Regression Model. Frequency and utilization as two parameters are selected in the model as cubic and primary, respectively. This is because, in the CMOS circuit model, frequency and power consumption relation are cubic. The second model is optimized based on Model 1. It added more terms with different exponents, with the highest order being 3. The reason for this modification is simply the intuitive idea that adding more terms will improve the accuracy of the model. We call it Model 2, the Multi-Term Regression Model. Considering the frequency/voltage levels, we further divide Model 2 into two different equations corresponding to 2 frequency/voltage levels. It uses one set of coefficients when the frequency is in the first 3 levels, and another set of coefficients when the frequency is in the later levels. We call it the Multi-Frequency Regression Model (Model 3). We improve Model 3 further and introduce Model 4, called the Per-Frequency Regression Model, which fits data to individual frequency.
Based on the test data from  III-B, we evaluated the performance of the four models and the results are summarized in the Table IV.
TABLE IV: The accuracy of regression models
@@
1 Simple Regression Model 0.2397 0.4184 -0.0242
2 Multi-Term Regression Model 0.0272 0.1365 0.8839
3 Multi-Frequency Regression Model 0.0270 0.1344 0.8848
4 Per-Frequency Regression Model 0.0182 0.1040 0.9221
@@
An $R^{2}$ score of 0.9221 implies that approximately 92.21% of the variance in power consumption can be explained by the model using frequency and utilization as predictors. This high $R^{2}$ value indicates a strong correlation between the predicted and actual power consumption values, signifying that the model fits the data well and has high predictive power.
The high $R^{2}$ score of the Per-Frequency Regression Model indicates that it can reliably predict the power consumption of the Jetson Nano Board (2GB) under various workloads and operating conditions.